# The communication complexity of distributed estimation

### v1

@@ 1.1 Motivation
Say that Alice holds a subset AA and Bob holds a subset $B$, of vertices in a graph $G=(V,E)$. Their goal is to estimate the density of edges between sets AA and $B$. Replacing each player’s input with the uniform distribution over their sets $A,B$, and using the function $f(i,j)=1[(i,j)\in E]$, the distributed estimation problem is to estimate the edge density $|E(A,B)|/(|A||B|)$.
@@ 1.3 Overview of lower bounds
The above results show that the variance reduction protocol in Theorem 2.1 is optimal in two important ways. First, the communication complexity of any protocol (irrespective of number of rounds) is at least $\Omega(n/\varepsilon)$ in general. Second, achieving the improved upper bound in Theorem 2.1 requires at least two rounds: 11-round protocols have a communication complexity of $\Omega(n/\varepsilon^{2})$.
@@ 2.1 Debiasing protocol for better variance reduction
Theorem $2.$
For any function $f:\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we have $\bar{\mathrm{R}}_{\varepsilon}(f)\leq\tilde{O}(\mathrm{R}(f)/\varepsilon)$.
@@
We will reduce from Gap Hamming on Inputs of length $2^{2k}$. We think of the inputs to Alice and Bob as matrices $A,B$ each in $\{\pm 1\}^{2^{k}\times 2^{k}}$. Let ${\langle A,B\rangle}_{F}$ denote the entry-wise or Frobenius inner product of AA and $B$. Computing this to within error $\varepsilon 2^{2k}$ requires $\Omega(\min(1/\varepsilon^{2},2^{2k}))$ bits of communication by Theorem 4.2.
@@ 2.2 Protocols based on spectral properties
For brevity, let $\sigma_{i}\equiv\sigma_{i}(f)$ for all ii. Alice and Bob first exchange their $\beta$-heavy elements $p^{h}$ and $q^{h}$. Then, for all $j\leq t$, they exchange $(p^{\ell})^{\mathsf{T}}u_{j}$ and $v_{j}^{\mathsf{T}}q^{\ell}$ to within precision $\varepsilon/\sigma_{1}$, which costs $\tilde{O}(t\log(\sigma_{1}/\varepsilon))$ bits. They then run Algorithm 3 to estimate the contribution from the singular vectors for $j>t$, which costs $\tilde{O}((\sigma_{t}/\varepsilon)^{2/3})$. ∎
@@ 3.1 $\mathsf{EQ}$
A simple 11-round protocol is where Alice sends Bob the list of her non-zero probability elements, along with the probabilities of those elements to within additive error $\varepsilon^{2}$. This uses $O(\log(1/\varepsilon))$ bits of communication per element. This requires $\tilde{O}(1/\varepsilon)$ bits of communication.
@@
Further, we can approximate each $A_{i}$ via a dyadic composition by a sum of multiples (with bounded coefficients) of $O(\log{\varepsilon}^{-1})$ matrices which are 11-sparse and have entries in $\{0,1\}$.
@@ 3.2 $\mathsf{GT}$
Note that given the interval partition, Bob can compute the sum (the second term in (5)) from just the values $p_{j}$. Additionally, for any specific value $x$, Bob can also compute the value inside the expectation in (5), since he knows $q_{j}$ and the distribution $q|_{j}$. Thus, Bob can estimate the value of the expectation given random samples from pp. This gives rise to the following protocol, illustrated in Algorithm 4: Alice and Bob share the partition, and Alice sends the values $p_{j}$ for each $j$. Alice also sends $k=O((\beta/\varepsilon)^{2})$ samples from pp. Bob then computes the sum in (5) and estimates the expectation using the samples.
@@
    Alice sends Bob $k=O((\beta/\varepsilon)^{2})$ samples from pp.
@@ 3.3 Absolute difference
    Alice sends Bob $k=O((\beta^{2}/\varepsilon)^{2})$ samples from pp.
@@
Let $f:[0,1]^{2}\to[-1,1]$ be a 11-Lipschitz convex function. Then, $\bar{\mathrm{R}}_{\varepsilon}(f)\leq\tilde{O}(1/\varepsilon^{2/5})$.
@@
Let $f:[0,1]\to[-1,1]$ be convex and 11-Lipschitz. Then there exist a probability measure $D$ on $[0,1]$ and a constant $c\in[-2,2]$ such that
@@
Consider the following communication problem: Alice has a convex function $f:[0,1]\to[-1,1]$ and Bob has a distribution $q$ on $[0,1]$, and their goal is to estimate $\operatorname*{\mathbf{E}}_{y\sim q}[f(y)]$ to within additive error $\varepsilon$. Note that this is an instance of the expectation estimation problem for absolute difference, since by 3.14, we can write $f$ as (a constant shift of) the expectation of |xy||x-y| as $x$ is drawn from some distribution pp. The work of [GORSS24] gives a one-way protocol for this problem with communication $\tilde{O}(1/\varepsilon^{2/3})$, by finding an approximate linear basis for the space of bounded convex functions.
@@ 4 A spectral technique for lower bounds
where the $\sigma_{i}$s are the singular values of AA.
If AA is invertible, then $1/\sigma_{i}$ are the singular values of $A^{-1}$ and $\lambda_{k}(A)$ is the trace norm of $A^{-1}$. By allowing $t<k$, we allow the flexibility of finding a subspace of the domain over which the inverse is well behaved.
@@
Given a matrix $A\in[-1,1]^{k\times k}$, and a function $f:\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we say $f$ realizes the matrix AA if there exist $\{x_{i}\in\mathcal{X}\}_{i\in[k]}$ and $\{y_{j}\in\mathcal{Y}\}_{j\in[k]}$ so that
@@
In the case where the domain is discrete, this is the same as saying that the communication matrix $M_{f}$ contains AA as a submatrix.
@@
Let $\{x_{i}\in\mathcal{X}\}_{i\in[k]}$ and $\{y_{j}\in\mathcal{Y}\}_{j\in[k]}$ be such that $A_{ij}=f(x_{i},y_{j})$. Let $r=rank(A)$, and let the singular-value-decomposition of AA be $A=U\Sigma V^{t}$ where we denote the columns of $U$ and $V$ by $\{u_{i}\in{\mathbb{R}}^{k}\}_{i\in[r]}$ and $\{v_{j}\in{\mathbb{R}}^{k}\}_{j\in[k]}$, and the singular values by $\{\sigma_{i}\}_{i\in[r]}$.
@@
To get a strong lower bound, we need to find matrices AA realized by $f$ where $\lambda_{t}(A)$ is as small as possible. As a sanity check, note that Theorem 2.2 guarantees that $t$ bits of communication suffice to achieve error $O(\mathrm{R}^{\mathrm{ow}}(f)/t)$. So the bound on $\varepsilon$ given by Theorem 4.5 cannot be asympototically stronger than $1/t$ (say $1/t^{0.9}$). Indeed, using the lower bound on $\lambda_{t}(A)$ from Equation 12,
@@ 4.1 Lower bounds for specific functions
The inner product function on $n$ bits realizes a Hadamard submatrix of size $2^{\ell}\times 2^{\ell}$ for every $\ell\leq n$. We have shown that $\lambda_{k}(H_{k})=\sqrt{k}$. This gives a lower bound of $\Omega(k)$ for $\varepsilon\leq 1/k$ for every $k$ that is a power of 22 bounded by $2^{n}$. This gives a $\Omega(1/\varepsilon)$ lower bound for every $\varepsilon\geq 2^{-n}$. ∎
@@
A similar lower bound applies to any function that realizes a $k\times k$ matrix AA where $\lambda_{k}(A)=O(k)$. In particular, we get that the protocol for $\mathsf{GT}_{n}$ from Theorem 3.7 is tight upto a logarithmic factor:
@@
The inverse of $T_{k}$ is the matrix $k\times k$ matrix $U_{k}$ where $U_{ii}=1$ and $U_{i,i+1}=-1$. It follows that $\|U_{k}\|\leq 2$, hence $\sigma_{i}(U_{k})\geq 1/\|U_{k}\|_{2}\geq 1/2$ for all ii. So $\lambda_{k}(\mathsf{GT}_{k})\leq 2k$. ∎
@@
$E_{k}^{-1}$ has $k$ entries bounded by 11 and $2k$ entries bounded by $1/2$, so
@@ 4.2 Lower bounds for Boolean functions with high rank
We may assume $\operatorname{rank}(f)\geq 4$, since otherwise we only need to show a lower bound of some constant, which holds trivially since $f$ is not constant. For $4\leq k\leq\operatorname{rank}(f)$, let $A\in\{0,1\}^{k\times k}$ be a full rank Boolean matrix realized by $f$. Since AA is Boolean, $\det(A)=1$. By Lemma 4.14,
@@ 5.1.1 Lower bound for computing $\mathsf{IP}_{n}^{t\times t}$
However, with a view towards extending to the case of general functions, we present an alternate, cleaner way. Note that in Section 5.1.1, we can further decompose $Z_{1}=Z_{1,B}+Z_{1,A}$. As the coordinates of $X$ in II are fixed, $Z_{1,B}$ only depends on $Y$; similarly $Z_{1,A}$ is a function of $X$ alone. Thus, we can condition on a typical value of $Z_{1,A},Z_{1,B}$ while mainting independence of $X,Y$.
@@ 5.1.2 Lower bound for computing $f^{t\times t}$ from discrepancy
Let us now return to the proof of Theorem 5.2, specifically the special case as alluded to in Claim $5.$
@@ 5.3 Discrepancy and max-divergence
(F3) Inclusion-maximal heavy implies immediate CBD. Fix $\rho>0$ and say that $(I,\alpha)$ is $\rho$-heavy for $\mu$ if $\log\big(\mu_{I}(\alpha)/\Theta^{I}(\alpha)\big)\ \geq\ \rho\,|I|$. If $(I,\alpha)$ is $\rho$-heavy and inclusion-maximal (no strict superset of II is $\rho$-heavy), then the conditional $\mu^{\prime}:=\mu\mid(X_{I}=\alpha)$ is $\rho$-CBD on the remaining coordinates: for every $U\subseteq S\setminus I$,
@@
Proof of (F3): Otherwise, for some $U$ we would have, by (23), $\mathrm{D}_{\infty}(\mu_{I\cup U}\|\Theta^{I\cup U})>\mathrm{D}_{\infty}(\mu^{\prime}_{U}\|\Theta^{U})+\log\frac{\mu_{I}(\alpha)}{\Theta^{I}(\alpha)}>\rho|U|+\rho|I|=\rho|I\cup U|,$ contradicting inclusion-maximality of II. ∎
@@
    Remove it: set $R_{i+1}:=R_{i}\setminus C_{i+1}$ and increment ii.
@@ 5.6 Expectations of pairwise products
If $y_{2}\notin\mathcal{G}_{Y}$, the inner absolute value is trivially at most 11. Hence, using (44) and (45),
@@ 6 One-way communication lower bound for parities
Choose the constant in the upper bound on NN so that $N\delta\leq 1/10$, hence $S\subseteq\widetilde{B}(S,y)$ w.h.p., i.e., no “misses.” Combining with (52) we get, with probability at least $4/5$ over the protocol coins,
@@
Given $m$ and the public coins, Alice can finish identifying $S$ to Bob by sending the index of $S$ as an NN-subset of $\widetilde{B}(S,y)$, using an additional
@@ Appendix A Omitted proofs
Since $f$ is convex and 11-Lipschitz on $[0,1]$, its right derivative $f^{\prime}_{+}(x)$ exists for every $x\in[0,1)$, is nondecreasing, right-continuous, and satisfies $f^{\prime}_{+}(x)\in[-1,1]$. Define $F:\mathbb{R}\to[0,1]$ by
@@
Both $f$ and $g$ are 11-Lipschitz on $[0,1]$, so $f^{\prime}$ and $g^{\prime}$ exist almost everywhere and equal $f^{\prime}_{+}$ and $g^{\prime}_{+}$ almost everywhere. It follows that $(f-g)^{\prime}(x)=0$ almost everywhere, so fgf-g is constant on $[0,1]$. Writing $c:=f(0)-g(0)=f(0)-\operatorname*{\mathbf{E}}[z]$ we obtain
@@
All derivatives of $\phi_{i}$ up to order i1i-1 (respectively of $\psi_{\ell-i}$ up to order $\ell-i-1$) vanish at 0 and 11. Integrating by parts $i$ times in $x$ and $\ell-i$ times in $y$ (there are no boundary terms since the derivatives vanish on the edge of the square) gives
@@
Here $A_{i,\ell}>0$ and $B_{i,\ell}<\infty$, both depending only on $\ell$ (and ii). Thus, picking $K_{\ell}$ to be larger than all ratios $B_{i,\ell}/A_{i,\ell}$ for $0\leq i\leq\ell$ yields a contradiction. Let $K_{\ell}\coloneqq\max_{0\leq i\leq\ell}\frac{B_{i,\ell}}{A_{i,\ell}}$. The previous display contradicts $m>K_{\ell}$. Therefore, in summary, we have

### v2

@@ 1.1 Motivation
Say that Alice holds a subset $A$ and Bob holds a subset $B$, of vertices in a graph $G=(V,E)$. Their goal is to estimate the density of edges between sets $A$ and $B$. Replacing each player’s input with the uniform distribution over their sets $A,B$, and using the function $f(i,j)=1[(i,j)\in E]$, the distributed estimation problem is to estimate the edge density $|E(A,B)|/(|A||B|)$.
@@ 1.3 Overview of lower bounds
The above results show that the variance reduction protocol in Theorem 2.1 is optimal in two important ways. First, the communication complexity of any protocol (irrespective of number of rounds) is at least $\Omega(n/\varepsilon)$ in general. Second, achieving the improved upper bound in Theorem 2.1 requires at least two rounds: $1$-round protocols have a communication complexity of $\Omega(n/\varepsilon^{2})$.
@@ 2.1 Debiasing protocol for better variance reduction
Theorem 2.2.
For any function $f:\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we have $\bar{\mathrm{R}}_{\varepsilon}(f)\leq\tilde{O}(\mathrm{R}^{\mathrm{ow}}(f)/\varepsilon)$.
@@
We will reduce from Gap Hamming on Inputs of length $2^{2k}$. We think of the inputs to Alice and Bob as matrices $A,B$ each in $\{\pm 1\}^{2^{k}\times 2^{k}}$. Let ${\langle A,B\rangle}_{F}$ denote the entry-wise or Frobenius inner product of $A$ and $B$. Computing this to within error $\varepsilon 2^{2k}$ requires $\Omega(\min(1/\varepsilon^{2},2^{2k}))$ bits of communication by Theorem 4.2.
@@ 2.2 Protocols based on spectral properties
For brevity, let $\sigma_{i}\equiv\sigma_{i}(f)$ for all $i$. Alice and Bob first exchange their $\beta$-heavy elements $p^{h}$ and $q^{h}$. Then, for all $j\leq t$, they exchange $(p^{\ell})^{\mathsf{T}}u_{j}$ and $v_{j}^{\mathsf{T}}q^{\ell}$ to within precision $\varepsilon/\sigma_{1}$, which costs $\tilde{O}(t\log(\sigma_{1}/\varepsilon))$ bits. They then run Algorithm 3 to estimate the contribution from the singular vectors for $j>t$, which costs $\tilde{O}((\sigma_{t}/\varepsilon)^{2/3})$. ∎
@@ 3.1 $\mathsf{EQ}$
A simple $1$-round protocol is where Alice sends Bob the list of her non-zero probability elements, along with the probabilities of those elements to within additive error $\varepsilon^{2}$. This uses $O(\log(1/\varepsilon))$ bits of communication per element. This requires $\tilde{O}(1/\varepsilon)$ bits of communication.
@@
Further, we can approximate each $A_{i}$ via a dyadic composition by a sum of multiples (with bounded coefficients) of $O(\log{\varepsilon}^{-1})$ matrices which are $1$-sparse and have entries in $\{0,1\}$.
@@ 3.2 $\mathsf{GT}$
Note that given the interval partition, Bob can compute the sum (the second term in (5)) from just the values $p_{j}$. Additionally, for any specific value $x$, Bob can also compute the value inside the expectation in (5), since he knows $q_{j}$ and the distribution $q|_{j}$. Thus, Bob can estimate the value of the expectation given random samples from $p$. This gives rise to the following protocol, illustrated in Algorithm 4: Alice and Bob share the partition, and Alice sends the values $p_{j}$ for each $j$. Alice also sends $k=O((\beta/\varepsilon)^{2})$ samples from $p$. Bob then computes the sum in (5) and estimates the expectation using the samples.
@@
    Alice sends Bob $k=O((\beta/\varepsilon)^{2})$ samples from $p$.
@@ 3.3 Absolute difference
    Alice sends Bob $k=O((\beta^{2}/\varepsilon)^{2})$ samples from $p$.
@@
Let $f:[0,1]^{2}\to[-1,1]$ be a $1$-Lipschitz convex function. Then, $\bar{\mathrm{R}}_{\varepsilon}(f)\leq\tilde{O}(1/\varepsilon^{2/5})$.
@@
Let $f:[0,1]\to[-1,1]$ be convex and $1$-Lipschitz. Then there exist a probability measure $D$ on $[0,1]$ and a constant $c\in[-2,2]$ such that
@@
Consider the following communication problem: Alice has a convex function $f:[0,1]\to[-1,1]$ and Bob has a distribution $q$ on $[0,1]$, and their goal is to estimate $\operatorname*{\mathbf{E}}_{y\sim q}[f(y)]$ to within additive error $\varepsilon$. Note that this is an instance of the expectation estimation problem for absolute difference, since by 3.14, we can write $f$ as (a constant shift of) the expectation of |xy||x-y| as $x$ is drawn from some distribution $p$. The work of [GORSS24] gives a one-way protocol for this problem with communication $\tilde{O}(1/\varepsilon^{2/3})$, by finding an approximate linear basis for the space of bounded convex functions.
@@ 4 A spectral technique for lower bounds
where the $\sigma_{i}$s are the singular values of $A$.
If $A$ is invertible, then $1/\sigma_{i}$ are the singular values of $A^{-1}$ and $\lambda_{k}(A)$ is the trace norm of $A^{-1}$. By allowing $t<k$, we allow the flexibility of finding a subspace of the domain over which the inverse is well behaved.
@@
Given a matrix $A\in[-1,1]^{k\times k}$, and a function $f:\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we say $f$ realizes the matrix $A$ if there exist $\{x_{i}\in\mathcal{X}\}_{i\in[k]}$ and $\{y_{j}\in\mathcal{Y}\}_{j\in[k]}$ so that
@@
In the case where the domain is discrete, this is the same as saying that the communication matrix $M_{f}$ contains $A$ as a submatrix.
@@
Let $\{x_{i}\in\mathcal{X}\}_{i\in[k]}$ and $\{y_{j}\in\mathcal{Y}\}_{j\in[k]}$ be such that $A_{ij}=f(x_{i},y_{j})$. Let $r=rank(A)$, and let the singular-value-decomposition of $A$ be $A=U\Sigma V^{t}$ where we denote the columns of $U$ and $V$ by $\{u_{i}\in{\mathbb{R}}^{k}\}_{i\in[r]}$ and $\{v_{j}\in{\mathbb{R}}^{k}\}_{j\in[k]}$, and the singular values by $\{\sigma_{i}\}_{i\in[r]}$.
@@
To get a strong lower bound, we need to find matrices $A$ realized by $f$ where $\lambda_{t}(A)$ is as small as possible. As a sanity check, note that Theorem 2.2 guarantees that $t$ bits of communication suffice to achieve error $O(\mathrm{R}^{\mathrm{ow}}(f)/t)$. So the bound on $\varepsilon$ given by Theorem 4.5 cannot be asympototically stronger than $1/t$ (say $1/t^{0.9}$). Indeed, using the lower bound on $\lambda_{t}(A)$ from Equation 12,
@@ 4.1 Lower bounds for specific functions
The inner product function on $n$ bits realizes a Hadamard submatrix of size $2^{\ell}\times 2^{\ell}$ for every $\ell\leq n$. We have shown that $\lambda_{k}(H_{k})=\sqrt{k}$. This gives a lower bound of $\Omega(k)$ for $\varepsilon\leq 1/k$ for every $k$ that is a power of $2$ bounded by $2^{n}$. This gives a $\Omega(1/\varepsilon)$ lower bound for every $\varepsilon\geq 2^{-n}$. ∎
@@
A similar lower bound applies to any function that realizes a $k\times k$ matrix $A$ where $\lambda_{k}(A)=O(k)$. In particular, we get that the protocol for $\mathsf{GT}_{n}$ from Theorem 3.7 is tight upto a logarithmic factor:
@@
The inverse of $T_{k}$ is the matrix $k\times k$ matrix $U_{k}$ where $U_{ii}=1$ and $U_{i,i+1}=-1$. It follows that $\|U_{k}\|\leq 2$, hence $\sigma_{i}(U_{k})\geq 1/\|U_{k}\|_{2}\geq 1/2$ for all $i$. So $\lambda_{k}(\mathsf{GT}_{k})\leq 2k$. ∎
@@
$E_{k}^{-1}$ has $k$ entries bounded by $1$ and $2k$ entries bounded by $1/2$, so
@@ 4.2 Lower bounds for Boolean functions with high rank
We may assume $\operatorname{rank}(f)\geq 4$, since otherwise we only need to show a lower bound of some constant, which holds trivially since $f$ is not constant. For $4\leq k\leq\operatorname{rank}(f)$, let $A\in\{0,1\}^{k\times k}$ be a full rank Boolean matrix realized by $f$. Since $A$ is Boolean, $\det(A)=1$. By Lemma 4.14,
@@ 5.1.1 Lower bound for computing $\mathsf{IP}_{n}^{t\times t}$
However, with a view towards extending to the case of general functions, we present an alternate, cleaner way. Note that in Section 5.1.1, we can further decompose $Z_{1}=Z_{1,B}+Z_{1,A}$. As the coordinates of $X$ in $I$ are fixed, $Z_{1,B}$ only depends on $Y$; similarly $Z_{1,A}$ is a function of $X$ alone. Thus, we can condition on a typical value of $Z_{1,A},Z_{1,B}$ while mainting independence of $X,Y$.
@@ 5.1.2 Lower bound for computing $f^{t\times t}$ from discrepancy
Let us now return to the proof of Theorem 5.2, specifically the special case as alluded to in Claim 5.5.
@@ 5.3 Discrepancy and max-divergence
(F3) Inclusion-maximal heavy implies immediate CBD. Fix $\rho>0$ and say that $(I,\alpha)$ is $\rho$-heavy for $\mu$ if $\log\big(\mu_{I}(\alpha)/\Theta^{I}(\alpha)\big)\ \geq\ \rho\,|I|$. If $(I,\alpha)$ is $\rho$-heavy and inclusion-maximal (no strict superset of $I$ is $\rho$-heavy), then the conditional $\mu^{\prime}:=\mu\mid(X_{I}=\alpha)$ is $\rho$-CBD on the remaining coordinates: for every $U\subseteq S\setminus I$,
@@
Proof of (F3): Otherwise, for some $U$ we would have, by (23), $\mathrm{D}_{\infty}(\mu_{I\cup U}\|\Theta^{I\cup U})>\mathrm{D}_{\infty}(\mu^{\prime}_{U}\|\Theta^{U})+\log\frac{\mu_{I}(\alpha)}{\Theta^{I}(\alpha)}>\rho|U|+\rho|I|=\rho|I\cup U|,$ contradicting inclusion-maximality of $I$. ∎
@@
    Remove it: set $R_{i+1}:=R_{i}\setminus C_{i+1}$ and increment $i$.
@@ 5.6 Expectations of pairwise products
If $y_{2}\notin\mathcal{G}_{Y}$, the inner absolute value is trivially at most $1$. Hence, using (44) and (45),
@@ 6 One-way communication lower bound for parities
Choose the constant in the upper bound on $N$ so that $N\delta\leq 1/10$, hence $S\subseteq\widetilde{B}(S,y)$ w.h.p., i.e., no “misses.” Combining with (52) we get, with probability at least $4/5$ over the protocol coins,
@@
Given $m$ and the public coins, Alice can finish identifying $S$ to Bob by sending the index of $S$ as an $N$-subset of $\widetilde{B}(S,y)$, using an additional
@@ Appendix A Omitted proofs
Since $f$ is convex and $1$-Lipschitz on $[0,1]$, its right derivative $f^{\prime}_{+}(x)$ exists for every $x\in[0,1)$, is nondecreasing, right-continuous, and satisfies $f^{\prime}_{+}(x)\in[-1,1]$. Define $F:\mathbb{R}\to[0,1]$ by
@@
Both $f$ and $g$ are $1$-Lipschitz on $[0,1]$, so $f^{\prime}$ and $g^{\prime}$ exist almost everywhere and equal $f^{\prime}_{+}$ and $g^{\prime}_{+}$ almost everywhere. It follows that $(f-g)^{\prime}(x)=0$ almost everywhere, so fgf-g is constant on $[0,1]$. Writing $c:=f(0)-g(0)=f(0)-\operatorname*{\mathbf{E}}[z]$ we obtain
@@
All derivatives of $\phi_{i}$ up to order i1i-1 (respectively of $\psi_{\ell-i}$ up to order $\ell-i-1$) vanish at 0 and $1$. Integrating by parts $i$ times in $x$ and $\ell-i$ times in $y$ (there are no boundary terms since the derivatives vanish on the edge of the square) gives
@@
Here $A_{i,\ell}>0$ and $B_{i,\ell}<\infty$, both depending only on $\ell$ (and $i$). Thus, picking $K_{\ell}$ to be larger than all ratios $B_{i,\ell}/A_{i,\ell}$ for $0\leq i\leq\ell$ yields a contradiction. Let $K_{\ell}\coloneqq\max_{0\leq i\leq\ell}\frac{B_{i,\ell}}{A_{i,\ell}}$. The previous display contradicts $m>K_{\ell}$. Therefore, in summary, we have